OpenStudy (leahhhmorgannn):

not getting any help in the Chemistry section and I really have no idea how to do this calculation?? How many moles of silver (Ag) are present in a sample of 3.8 × 1024 atoms Ag? 1.6 mol Ag 2.3 mol Ag 6.3 mol Ag 22.8 mol Ag

OpenStudy (anonymous):

To convert from atoms to moles, simply divide the atom amount by Avogadro's number. Like so: \[\frac{ 3.8 \times 10^{24} }{ 6.022 \times 10^{23} } ~~~Ag\] And you'll get your answer.

OpenStudy (seeing.in.pixels):

@leahhhmorgannn asking for help in biology won't always work. You see, chemistry usually comes after biology, so people in here usually don't know how to help. It worked our this time, though, so it's all good :) - Pixels

OpenStudy (leahhhmorgannn):

@Seeing.In.Pixels I had asked in Chemistry around 10 to 20 minutes prior to asking in here. I was getting no help, no matter who I tagged so I tried here because I have some friends in this section from last year. I get that Chem comes after Biology - I took Biology last year. But I wasn't initially getting help here either, so I finally got someone to help explain it to me in the Math section.

OpenStudy (seeing.in.pixels):

Oh, okay. I'm sorry. It was more of a suggestion. Still, you're right, and I apologize. Love, Pixels